OpenStudy (anonymous):

Are the graphs of the lines in the pair parallel? Explain. y = 6x + 9 27x – 3y = –81 A. No, since the slopes are different. B. No, since the y-intercepts are different. C. Yes, since the slopes are the same and the y-intercepts are the same. D. Yes, since the slopes are the same and the y-intercepts are different.

OpenStudy (anonymous):

27x-3y=-81 -27x -27x -3y=-81-27x /-3 /-3 /-3 y=27+9x y=9x+27

OpenStudy (anonymous):

both the slopes and y intercepts are different so we can conclude it is neither c nor d

OpenStudy (anonymous):

if two lines have the same y intercept they can not be parallel because that would mean they are crossing each other

OpenStudy (anonymous):

the slopes are the same however and the y intercepts are different, they are parallel

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

try typing in your equation here next time to see